OpenStudy (amistre64):

\[2+\cfrac{1}{2+\cfrac{1}{4+\cfrac{1}{2+1\cfrac{4+...}}}}\]

OpenStudy (amistre64):

that dint work out so well

OpenStudy (anonymous):

not really, can you attach a picture?

OpenStudy (amistre64):

\[2+\cfrac{1}{2+\cfrac{1}{4+\cfrac{1}{2+\cfrac{1}{4+\frac{1}{...}}}}}\]

OpenStudy (amistre64):

the continued fraction expansion of sqrt(6) :)

OpenStudy (amistre64):

\[\Large\begin{array}l \int_{0}^{b}x\ cos(x)-x^{10}dx\\\\ =cos(x)-x\ sin(x)-10x^{9}; [0,b]\\\\ =cos(b)-b\ sin(b)-10b^{9}-cos(0)\\\\ =cos(b)-b\ sin(b)-10b^{9}-1\\\\ \end{array} \]

OpenStudy (amistre64):

\[\Large\begin{array}l \int_{0}^{b}x\ cos(x)-x^{10}dx\\\\ =x\ sin(x)+cos(x)-\frac{x^{11}}{11}; [0,b]\\\\ =b\ sin(b)+cos(b)-\frac{b^{11}}{11} -cos(0)\\\\ =b\ sin(b)+cos(b)-\frac{b^{11}}{11} -1\\\\ \end{array}\]

OpenStudy (amistre64):

\[\Large\begin{array}l \int_{0}^{\sqrt{3-\sqrt{6}}}x^5-6x^3+4x-x\ dx\\\\ +\int_{\sqrt{3-\sqrt{6}}}^{\sqrt{3+\sqrt{6}}}x^5-6x^3+4x-x\ dx\\\\ A=\frac{x^6}{6}-\frac{3x^4}{2}+\frac{3x^2}{2}\\\\ =\frac{(\sqrt{3-\sqrt{6}})^6}{6}-\frac{3(\sqrt{3-\sqrt{6}})^4}{2}+\frac{3(\sqrt{3-\sqrt{6}})^2}{2}\\\\ =\frac{(81-33\sqrt{6})}{6}-\frac{3(15-6\sqrt{6})}{2}+\frac{3(3-\sqrt{6})}{2}\\\\ =\frac{(81-33\sqrt{6})-9(15-6\sqrt{6})+9(3-\sqrt{6})}{6}\\\\ =\frac{81-33\sqrt{6}-135+54\sqrt{6}+27-9\sqrt{6}}{6}\\\\ =\frac{-27+12\sqrt{6}}{6}\\\\ =\frac{-9+4\sqrt{6}}{2}\\\\ +\int_{\sqrt{3-\sqrt{6}}}^{\sqrt{3+\sqrt{6}}}-(x^5-6x^3+4x)+x\ dx\\\\ B=-\frac{x^6}{6}+\frac{3x^4}{2}-\frac{3x^2}{2}\\\\ =-\frac{x^6}{6}+\frac{3x^4}{2}-\frac{3x^2}{2}+A\\\\ =-\frac{(\sqrt{3+\sqrt{6}})^6}{6}+\frac{3(\sqrt{3+\sqrt{6}})^4}{2}-\frac{3(\sqrt{3+\sqrt{6}})^2}{2}\\\\+A\\\\ =\frac{-(\sqrt{3+\sqrt{6}})^6+9(\sqrt{3+\sqrt{6}})^4-9(\sqrt{3+\sqrt{6}})^2}{6}\\\\+A\\\\ =\frac{ -(81+33\sqrt{6})+9(15+6\sqrt{6})-9(3+\sqrt{6})}{6}\\\\+A\\\\ =\frac{ -81-33\sqrt{6}+135+54\sqrt{6}-27-9\sqrt{6}}{6}+A\\\\ =\frac{ 9+4\sqrt{6}}{2}+A\\\\ =\frac{ 9+4\sqrt{6}-9+4\sqrt{6}}{2}\\\\ =\frac{8\sqrt{6}}{2}\\\\ A+B=\frac{-9+4\sqrt{6}}{2}+\frac{8\sqrt{6}}{2}\\\\ =\frac{-9+12\sqrt{6}}{2}\\\\ 12\sqrt{6}-9 \end{array}\]